Need help with the following accounting problem.

Required information Use the following information for the Problems below. [The following information applies to the questionDuring the current quarter, the company operated at 90% of capacity and produced 54,900 units of product; actual direct laborProblem 23-4A Computation of materials, labor, and overhead variances LO P2, P3 Required: 1. Compute the direct materials cos2. Compute the direct labor cost variance, including its rate and efficiency variances. Actual Cost Standard Cost3. Compute the overhead controllable and volume variances. Controllable Variance Actual overhead Budgeted overhead Controllab
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1364000 $7.80 $10,639,200.00 1364000 $5.00 $6,820,000.00 1647000 $5.00 $8,235,000.00
$3,819,200.00 Unfavorable $1,415,000.00 Favorable
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $3,819,200.00 Unfavorable
Direct material quantity variance $1,415,000.00 Favorable
Total direct material variance $2,404,200.00 Unfavorable

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
316300 $11.10 $3,510,930.00 316300 $14.00 $4,428,200.00 384300 $14.00 $5,380,200.00
$917,270.00 Favorable $952,000.00 Favorable
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $917,270.00 Favorable
Direct Labor Efficiency variance $952,000.00 Favorable
Total direct labor variance $1,869,270.00 Favorable

Solution 3:

Controllable Variance
Actual overhead $5,090,600.00
Budgeted overhead $5,764,500.00
Controllable variance $673,900.00 Favorable
Fixed overhead volume variance
Budgeted fixed overhead $3,074,400.00
Fixed overhead cost applied $3,458,700.00
Fixed overhead volume variance $384,300.00 Favorable